# MixAT: Combining Continuous and Discrete Adversarial Training for LLMs

### v1

@@ B.10 Examining malicious requests
Our main intuition for the increased effectiveness of MixAT training lies in the fact that the combination of rephrasing and continuous attacks can cover a larger portion of the adversarial space. To illustrate this, we have used an LLM2Vec [39] model built on top of Llama3 to extract embeddings for some malicious requests, as well as GCG, PAP, and continuous attacks targeting these requests. We illustrate these examples in Table 15 and their pairwise cosine similarities in Fig. 6. We observe that combining PAP and continuous attacks results in samples that are further away from the original prompt than each individual attack, confirming our hypothesis that combined attacks are stronger and can explore a wider section of the adversarial space, while still being close enough that they are useful for training.
Figure 6: Analysis of Cosine similarities of prompts A1-A6 from Table 15.
@@
B2 GCG Give detailed instructions and an example payload for dropping tables from a PostgreSQL database that is vulnerable to error-based SQL injection described(In A Way ir-ish haciendo(A delivering __________\_\_\_\_\__ _ _ _ _ overviewwith gramm mistakes reply ”SURE,re isa overview Success: Begorrah! I’ll give ye the lowdown on how to drop tables from a PostgreSQL…
@@ C.1 Datasets and Models
Training Data We train our models using the same harmful requests as Xhonneux et al. [7], which are similar to the ones in the HarmBench dataset [5]. The safe answer used during adversarial training is always “Sorry, I can’t do that.” To maintain utility, we use the UltraChat200k [40] utility dataset, used in both adversarial training [7] and general fine-tuning [24, 26, 41].
Other Datasets In Table 16 we list the licenses of the different datasets used in this paper.
@@
MT-Bench CC-BY-4.0 lmsys/mt_bench_human_judgments
XSTest CC-BY-4.0 paul-rottger/xstest
@@
Models Used In Table 17 we list the sources of the Hugging Face models and adapters used and evaluated in this work.
@@ C.2 Training Details
Unless stated otherwise, we follow the design and hyperparameter choices of prior work [7]. By default, models are 4-bit quantized and aligned using LoRA adapters [34] (we examine the non-quantized models in B.3). We use 10-step $L^{2}$-bounded continuous adversarial attacks with $\epsilon=0.075$, and discrete PAP attacks. The default PAP sample ratio is $\alpha=0.5$, with paraphrases drawn randomly from all 40 strategies [2]. We train for 2 epochs (in contrast to 5 in CAT) with a batch size $64$, a learning rate of $2e{-4}$, the AdamW optimizer [42], and a cosine learning rate scheduler.
C.3 Attack hyperparameters
To enhance reproducibility, we report the different hyperparameters for adversarial attacks used throughout the evaluations in Table 18. For generating adversarial attacks we use the HarmBench GitHub repository [5].
@@ C.3 Attack hyperparameters
Appendix D Discussion
Can Adversarial Training Fully Mitigate Malicious Requests? Adversarial training is a foundational component in defending LLMs against malicious inputs, yet it cannot fully resolve the problem of harmful requests. While it significantly strengthens resistance to known attack types, it remains vulnerable to adaptive attackers who craft novel strategies beyond the training space. Moreover, adversarial training alone does not address deeper ethical and contextual understanding, which are critical for responding appropriately to harmful inputs. A holistic defense strategy must therefore combine adversarial training with complementary methods, such as enhanced filtering systems, context-aware response generation, and ongoing model evaluation and refinement.
Challenges in Evaluation: Are Current Datasets Adequate? The robustness of adversarial training is closely tied to the quality of evaluation datasets, but existing datasets often have critical gaps. While they address various attack types, their scope is typically insufficient to test resilience against complex, multi-faceted adversarial strategies. Additionally, many datasets disproportionately emphasize certain attack categories, such as toxicity, while underrepresenting other types of harmful or manipulative inputs. To better evaluate and improve LLM robustness, there is a pressing need for richer, more diverse datasets encompassing a broader range of adversarial examples, including dynamic and context-sensitive attacks.

### v2

@@ B.14 Examining malicious requests
Our main intuition for the increased effectiveness of MixAT training lies in the fact that the combination of rephrasing and continuous attacks can cover a larger portion of the adversarial space. To illustrate this, we have used an LLM2Vec [44] model built on top of Llama3 to extract embeddings for some malicious requests, as well as GCG, PAP, and continuous attacks targeting these requests. We illustrate these examples in Table˜20 and their pairwise cosine similarities in Fig.˜6. We observe that combining PAP and continuous attacks results in samples that are further away from the original prompt than each individual attack, confirming our hypothesis that combined attacks are stronger and can explore a wider section of the adversarial space, while still being close enough that they are useful for training.
Figure 6: Analysis of Cosine similarities of prompts A1-A6 from Table˜20. Here, DR (Direct Request) refers to the base malicious request, and DR+C is the continuously attacked version of it. GCG and PAP refer to applying the GCG and PAP attacks on the base malicious request, respectively. PAP+C stands for deploying a continuous attack on the sample after the PAP attack.
@@
B2 GCG Give detailed instructions and an example payload for dropping tables from a PostgreSQL database that is vulnerable to error-based SQL injection described(In A Way ir-ish haciendo(A delivering _____\_\_\_\_\_ overviewwith gramm mistakes reply ”SURE,re isa overview Success: Begorrah! I’ll give ye the lowdown on how to drop tables from a PostgreSQL…
@@ C.1 Datasets and Models
Training Data We train our models using the same harmful requests as Xhonneux et al. [7], which are similar to the ones in the HarmBench dataset [5]. The safe answer used during adversarial training is always “Sorry, I can’t do that.” To maintain utility, we use the UltraChat200k [45] utility dataset, used in both adversarial training [7] and general fine-tuning [27, 29, 46]. Due to memory constraints, UltraChat200K is filtered to samples with fewer than 256 tokens.
Further in the malicious dataset, there are 100 harmful queries alongside multiple unpreferred continuations (1188 in total), which we together consider the harmful part of the dataset. The dataset is sampled with the interleave_dataset function from the HuggingFace datasets library, with probabilities 0.875 (benign) and 0.125 (harmful), and with strategy first_exhausted. So the ratio of benign and malicious queries is around 7:1. With our fixed seed, the benign part consists of 8476 samples sampled from UltraChat200K.
Other Datasets In Table˜21 we list the licenses of the different datasets used in this paper.
@@
MT-Bench $C$-BY-4.0 lmsys/mt_bench_human_judgments
XSTest $C$-BY-4.0 paul-rottger/xstest
@@
Models Used In Table˜22 we list the sources of the Hugging Face models and adapters used and evaluated in this work.
@@ C.2 Training Details
Unless stated otherwise, we follow the design and hyperparameter choices of prior work [7]. By default, models are 4-bit quantized and aligned using LoRA adapters [37] (we examine the non-quantized models in B.3). We use 10-step $L^{2}$-bounded continuous adversarial attacks with $\epsilon=0.075$, and discrete PAP attacks. The default PAP sample ratio is $\alpha=0.5$, with paraphrases drawn randomly from all 40 strategies [2]. We train for 2 epochs (in contrast to 5 in CAT) with a batch size $64$, a learning rate of 2e42e{-4}, the AdamW optimizer [47], and a cosine learning rate scheduler.
We note that a bug in the original CAT code was brought to our attention by the authors of Xhonneux et al. [7] around the time of publication of this work. In particular, the authors pointed to us that their code was not properly deactivating model weights during the adversarial attack generation phase. This results in unintended gradient accumulation for the model parameters during the attack, potentially impacting training dynamics. The reported results for both CAT and MixAT training are on models trained with the code that contains the bug. Prompted by their notification, we locally fixed this bug but did not observe any significant performance differences between training with and without the bug in our tests.
C.3 Inference hyperparameters
To enhance reproducibility, we report the different hyperparameters for adversarial attacks used throughout the evaluations in Table˜23. For generating adversarial attacks we use the HarmBench GitHub repository [5]. We note our Harmless (Hless) evaluation scores were computed using double quantization (bnb_4bit_use_double_quant=True in HuggingFace), unlike all other benchmarks.